# Vim-F: Visual State Space Model Benefiting from Learning in the Frequency Domain

### v2

@@ Abstract
In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as the Mamba deep learning models, have made significant progress in modeling long sequences. Compared to traditional convolutional neural networks (CNNs) and Vision Transformers (ViTs), the performance of Vision Mamba (ViM) methods is not yet fully competitive. To enable SSMs to process image data, ViMs typically flatten 2D images into 1D sequences, inevitably ignoring some 2D local dependencies, thereby weakening the model’s ability to interpret spatial relationships from a global perspective. We believe that the introduction of frequency domain information can enable ViM to achieve a better global receptive field during the scanning process. We propose a novel model called Vim-F, which employs pure Mamba encoders and scans in both the frequency and spatial domains. Moreover, we question the necessity of position embedding in ViM and remove it accordingly in Vim-F, which helps to fully utilize the efficient long-sequence modeling capability of ViM. Finally, we redesign a patch embedding for Vim-F, leveraging a convolutional stem to capture more local correlations, further improving the performance of Vim-F. Code is available at: https://github.com/yws-wxs/Vim-F.
Index Terms:
Vision Mamba, State Space Model, Fast Fourier Transform, Scanning strategy
I Introduction
@@
Visual representation learning stands as a pivotal research topic within the realm of computer vision. In the early era of deep learning, Convolutional Neural Networks (CNNs) occupied a dominant position. CNNs have several built-in inductive biases that make them well-suited to a wide variety of computer vision applications. In addition, CNNs are inherently efficient because when used in a sliding-window manner, the computations are shared. Inspired by the significant success of the Transformer architecture in the field of Natural Language Processing (NLP), researchers have recently applied Transformer to computer vision tasks. Except for the initial ”patch embedding” module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible. Compared to CNNs, ViTs typically exhibit superior performance, which can be attributed to the global receptive field and dynamic weights facilitated by the attention mechanism. However, the attention mechanism requires quadratic complexity in terms of image sizes, resulting in significant computational cost when addressing downstream dense prediction tasks, such as object detection and instance segmentation.
Figure 1: The overview of the Vim model.
@@
Derived from the classic Kalman filter model, modern State Space Models (SSMs) excel at capturing long-range dependencies and reap the benefits of parallel training. The Visual Mamba Models (ViMs) [1, 2, 3, 4], which are inspired by recently proposed SSMs [5, 6], utilize the Selective Scan Space State Sequential Model (S6) to compress previously scanned information into hidden states, effectively reducing quadratic complexity to linear. Most of these studies integrate the original SSM structure of Mamba into their basic models, thereby necessitating the conversion of 2D images into 1D sequences for SSM-based processing. However, due to the non-causal nature of visual data, flattening spatial data into one-dimensional tokens destroys local two-dimensional dependencies, impairing the model’s capacity to accurately interpret spatial relationships. Vim [1] addresses this issue by scanning in bidirectional horizontal directions, while VMamba [2] adds vertical scanning, enabling each element in the feature map to integrate information from other locations in different directions. Subsequent works, such as LocalMamba [3] and EfficientVMamba [4], have designed a series of novel scanning strategies in the spatial domain, aimed at capturing local dependencies while maintaining a global view. We believe that these strategies do not significantly enhance the model’s ability to obtain a better global receptive field. These efforts have prompted us to focus on the value of scanning in the frequency domain. We know that the use of 2D Discrete Fourier transform (DFT) to convert feature maps into spectrograms does not alter their shapes. The value $P(x,y)$ at any point $(x,y)$ in the spectrogram depends on the entire original feature map. Therefore, scanning in the frequency domain ensures that the model always has a good global receptive field. On the other hand, without considering spectral shifts, the high-frequency components are located in the corners of the spectrogram, whereas the low-frequency components are centered. Scanning across the spectrogram typically alternates between accessing low and high frequencies, which may facilitate balanced modeling. Therefore, we adopt the most direct solution, which is to add the amplitude spectrum of the feature map to the original feature map, allowing the visual Mamba encoder to learn the fused semantic information. Using the Fast Fourier Transform (FFT) algorithm to calculate the amplitude spectrum not only introduces no trainable parameters but also has minimal impact on performance. To clearly demonstrate the effectiveness and ease of use of our work, we chose Vim as the base model, which to our knowledge is the first pure-SSM-based visual model. Based on the open-source code provided by the authors, our method, called Vim-F, can be easily implemented by adding a few lines of code. Secondly, we improve the patch embedding to capture more local dependencies. Many successful lightweight transformer models, such as EfficientFormer [7] and MobileViT [8], start their networks with a convolutional stem as patch embedding. This is attributed to the inherent spatial inductive bias of CNN, which is beneficial for capturing local information. In addition, CNNs are less sensitive to data augmentation, and the training process of a hybrid architecture is more stable. Inspired by these works, we design a patch embedding for Vim-F. Specifically, we use multiple convolutional layers for downsampling and then flatten the feature maps. Finally, we argue that the Mamba model is essentially still a recurrent neural network (RNN). Therefore, Vim lacks the rationale for using positional embeddings, and thus our Vim-F does not employ positional embeddings. Additionally, we reduce the model’s sensitivity to the order of input sequences by alternately stacking linear attention blocks and Vim-F blocks, while maintaining overall linear complexity. We refer to this hybrid architecture as Vim-F(H). We outline our contributions below:
  • Based on Vim [1], we propose Vim-F, which enables SSM to scan both in the spatial and frequency domains, ensuring that SSM always has a global view during the scanning process, thereby enhancing the model’s ability to interpret spatial relationships. To the best of our knowledge, our work is the first to propose scanning in the frequency domain.
  • We design a novel patch embedding module that utilizes a mixture of overlapping and non-overlapping convolutions with small kernel sizes to downsample the feature map and then flatten it. Additionally, we question the necessity of position embedding in Vim, thus Vim-F does not utilize position embedding.
  • The experimental results demonstrate that the simple Vim variant, Vim-F, achieves better performance while maintaining similar computational costs and parameters. Additionally, introducing linear attention to reduce Vim-F’s sensitivity to the order of input sequences can further enhance the model’s performance.
II Related Works
II-A State Space Models
@@
In mathematics, State Space Models (SSMs) are usually expressed as linear ordinary differential equations (ODEs). These models transform input $D$-dimensional sequence $x(t)\in\mathbb{R}^{L\times D}$ into output sequence $y(t)\in\mathbb{R}^{L\times D}$ by utilizing a learnable latent state $h(t)\in\mathbb{R}^{L\times N}$ that is not directly observable. The mapping process could be denoted as:
@@
II-B Overview of Vim
As mentioned earlier, the effectiveness of our work is demonstrated through the improvement of Vim. Therefore, we provide a brief introduction to the overall architecture of Vim. As shown in Fig. 1, Vim first reshapes the 2D image $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{n\times(p^{2}\cdot c)}$, where $c$ is the number of channels, $(h,w)$ is the resolution of the original image, and $(p,p)$ is the resolution of each image patch. The effective sequence length for Vim is therefore $n=hw/p^{2}$. Then, Vim linearly projects $\mathbf{X}_{p}$ into vectors of size $d$ and adds them to the position embeddings $E_{pos}\in\mathbb{R}^{(n+1)\times d}$, as follows:
@@ II-B Overview of Vim
where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of blocks, and $\mathbf{Norm}$ is the normalization layer.
Figure 2: Illustrating the scanning strategy of Vim-F through an example of forward scanning in both the frequency and spatial domains. Due to the translation invariance of the Fourier transform, the Fourier transforms of figures (a) and (b) are identical.
III Rethinking the Scanning Strategy of ViMs
@@
The non-causal nature of two-dimensional spatial patterns in images is inherently inconsistent with the causal processing framework of SSM. Therefore, it is meaningful to explore methods for serializing images. However, the majority of existing work on scanning strategies is typically based on a rather idealized assumption that introducing more or more complex inductive biases regarding the order of one-dimensional sequences can potentially enhance Mamba’s ability to understand local image correlations. Theoretically, regardless of the order, reducing the distance between adjacent tokens in one direction in two-dimensional space inevitably increases the distance between adjacent tokens in another direction, thus these strategies are essentially equivalent. Additionally, in the absence of global information serving as a ”frame of reference,” these sorting methods are essentially disconnected, making it inherently difficult for Mamba to gain additional spatial relationship information by scanning multiple sequences. In practice, works involving scanning strategy research, including Vim [1], VMamba [2], LocalMamba [3], EfficientVMamba [4], and PlainMamba [9], lack clear comparative experiments to verify the necessity and effectiveness of their scanning strategies. Zhu et al. [10] conduct a comprehensive experimental investigation on the impact of mainstream scanning directions and their combinations on semantic segmentation of remotely sensed images. Through extensive experiments on the LoveDA, ISPRS Potsdam, and ISPRS Vaihingen datasets, they demonstrate that no single scanning strategy outperforms others, regardless of their complexity or the number of scanning directions involved. In summary, an effective scanning strategy needs to maintain a global receptive field throughout the scanning process and demonstrate its effectiveness through analysis and comparative experiments.
IV Methodology
@@
In this section, we first elaborate on the significance and benefits of frequency-domain scanning. Subsequently, we present the specific details of patch embedding in Vim-F. Finally, we discuss the composition of the Vim-F and the Vim-F(H) block.
IV-A 2D Scanning in the Frequency Domain
The flattening method used in Vim [1] disrupts these local dependencies and significantly increases the distance between vertically adjacent tokens. We address this limitation by introducing frequency-domain scanning. Specifically, we first perform a Fourier transform on the feature map to obtain:
$F(u,v)=\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}f(x,y)e^{-j2\pi\left(\frac{ux}{H}+\frac{vy}{W}\right)},$ (4)
where $j$ represents the imaginary unit, $H$ and $W$ represent the height and width of the feature map, respectively. $f(x,y)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(x,y)$ in the feature map, and $F(u,v)$ represents the corresponding value at the coordinate $(u,v)$ in the frequency domain. $|F(u,v)|\in\mathbb{R}^{H\times W}$ represents the modulus of $F(u,v)$, which is also known as the amplitude spectrum. Fourier transform has translation invariance. In other words, if a function is translated in the spatial domain, its Fourier transform remains the same as the original function. This property is quite meaningful for addressing the issue of increasing distances between vertically adjacent tokens, as shown in Fig. 2. The amplitude spectra of figures (a) and (b) are identical, so scanning in the frequency domain helps reduce the inductive bias introduced by the scanning strategy. Therefore, frequency domain scanning does not require fancy strategies, and it can be consistent with Vim’s original spatial domain scanning strategy.
IV-B Patch Embedding

### v3

@@ Abstract
In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as Mamba deep learning models, have made significant progress in modeling long sequences. Compared to Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), Vision Mamba (ViM) methods have not yet achieved fully competitive performance. To enable SSMs to process image data, ViMs typically flatten 2D images into 1D sequences, inevitably ignoring some 2D local dependencies, thereby weakening the model’s ability to interpret spatial relationships from a global perspective. We believe that the introduction of frequency domain information can enable ViM to achieve a better global receptive field during the scanning process. We propose a novel model called Vim-F, which employs pure Mamba encoders and scans in both the frequency and spatial domains. Moreover, considering that Mamba remains essentially a Recurrent Neural Network (RNN), we question the necessity of position embedding in ViM and remove it accordingly in Vim-F. Vim-F has good scalability. As far as we know, its variant Vim-F(CF) is the first ViM model to use a convolution-free ViM encoder. Another variant, Vim-F(H), introduces a linear attention mechanism. This reduces the model’s sensitivity to the input sequence and achieves better performance.
1 Introduction
@@
Visual representation learning stands as a pivotal research topic within the realm of computer vision. In the early era of deep learning, Convolutional Neural Networks (CNNs) occupied a dominant position. CNNs have several built-in inductive biases that make them well-suited to a wide variety of computer vision applications. In addition, CNNs are inherently efficient because when used in a sliding-window manner, the computations are shared. Inspired by the significant success of the Transformer architecture in the field of Natural Language Processing (NLP), researchers have recently applied Transformer to computer vision tasks. Except for the initial “patch embedding” module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible. Compared to CNNs, ViTs typically exhibit superior performance, which can be attributed to the global receptive field and dynamic weights facilitated by the attention mechanism. However, the attention mechanism requires quadratic complexity in terms of image sizes, resulting in significant computational cost when addressing downstream dense prediction tasks, such as object detection and instance segmentation.
@@
Derived from the classic Kalman filter model, modern State Space Models (SSMs) excel at capturing long-range dependencies and reap the benefits of parallel training. The Visual Mamba Models (ViMs) [22, 9, 5, 14], which are inspired by recently proposed SSMs [2, 12], utilize the Selective Scan Space State Sequential Model (S6) to compress previously scanned information into hidden states, effectively reducing quadratic complexity to linear. Most of these studies integrate the original SSM structure of Mamba into their basic models, thereby necessitating the conversion of 2D images into 1D sequences for SSM-based processing. However, due to the non-causal nature of visual data, flattening spatial data into one-dimensional tokens destroys local two-dimensional dependencies, impairing the model’s capacity to accurately interpret spatial relationships. Vim [22] addresses this issue by scanning in bidirectional horizontal directions, while VMamba [9] adds vertical scanning, enabling each element in the feature map to integrate information from other locations in different directions. Subsequent works, such as LocalMamba [5] and EfficientVMamba [14], have designed a series of novel scanning strategies in the spatial domain, aimed at capturing local dependencies while maintaining a global view. We believe that these strategies do not significantly enhance the model’s ability to obtain a better global receptive field. These efforts have prompted us to focus on the value of scanning in the frequency domain.
We know that the use of 2D Discrete Fourier transform (DFT) to convert feature maps into spectrograms does not alter their shapes. The value $P(x,y)$ at any point $(x,y)$ in the spectrogram depends on the entire original feature map. Therefore, scanning in the frequency domain ensures that the model always has a good global receptive field. On the other hand, without considering spectral shifts, the high-frequency components are located in the corners of the spectrogram, whereas the low-frequency components are centered. Scanning across the spectrogram typically alternates between accessing low and high frequencies, which may facilitate balanced modeling. Therefore, we adopt the most direct solution, which is to add the amplitude spectrum of the feature map to the original feature map, allowing the visual Mamba encoder to learn the fused semantic information. Using the Fast Fourier Transform (FFT) algorithm to calculate the amplitude spectrum not only introduces no trainable parameters but also has minimal impact on performance. To clearly demonstrate the effectiveness and ease of use of our work, we chose Vim as the base model, which to our knowledge is the first pure-SSM-based visual model. Based on the open-source code provided by the authors, our method, called Vim-F, can be easily implemented by adding a few lines of code. Secondly, we improve the patch embedding to capture more local dependencies. Many successful lightweight transformer models, such as EfficientFormer [8] and MobileViT [13], start their networks with a convolutional stem as patch embedding. This is attributed to the inherent spatial inductive bias of CNN, which is beneficial for capturing local information. In addition, CNNs are less sensitive to data augmentation, and the training process of a hybrid architecture is more stable. Inspired by these works, we design a patch embedding for Vim-F. Specifically, we use multiple convolutional layers for downsampling and then flatten the feature maps. Finally, we argue that the Mamba model is essentially still a recurrent neural network (RNN). Therefore, Vim lacks the rationale for using positional embeddings, and thus our Vim-F does not employ positional embeddings. Additionally, we reduce the model’s sensitivity to the order of input sequences by alternately stacking linear attention blocks and Vim-F blocks, while maintaining approximately linear complexity. We refer to this hybrid architecture as Vim-F(H). Finally, we explore the significance and feasibility of the convolution-free visual Mamba encoder and propose a proof-of-concept network, Vim-F(CF). Through extensive experiments on image classification, object detection, instance segmentation, and semantic segmentation tasks, the simple Vim variants, Vim-F(H) and Vim-F(CF), achieve better performance.
2 Related Works
@@
State Space Models. In mathematics, State Space Models (SSMs) are usually expressed as linear ordinary differential equations (ODEs). These models transform input $D$-dimensional sequence $x(t)\in\mathbb{R}^{L\times D}$ into output sequence $y(t)\in\mathbb{R}^{L\times D}$ by utilizing a learnable latent state $h(t)\in\mathbb{R}^{L\times N}$ that is not directly observable. The mapping process could be denoted as:
@@
Discretization. The primary objective of discretization is to convert ODEs into discrete functions, allowing the model to align with the sampling frequency of the input signal for more efficient computation. Following the work [3], the continuous parameters ($\bm{A}$, $\bm{B}$) can be discretized by zero-order hold rule with a given sample timescale $\Delta\in\mathbb{R}^{N}$:
$\bar{\bm{A}}=e^{\Delta\bm{A}},$ (2)
$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(e^{\Delta\bm{A}}-\bm{I})\cdot\Delta\bm{B},$
$\bar{\bm{C}}=\bm{C},$
$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$
$y(t)=\bar{\bm{C}}h(t),$
where $\bm{\bar{A}}\in\mathbb{R}^{N\times N}$, $\bm{\bar{B}}\in\mathbb{R}^{D\times N}$ and $\bm{\bar{C}}\in\mathbb{R}^{N\times D}$.
To improve computational efficiency, the iterative process described in equation 2 can be accelerated through parallel computing using global convolutional operations:
$\begin{split}\bm{y}&=\bm{x}\circledast\bm{\overline{K}}\\ \text{with}\quad\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{L-1}\bm{\overline{B}}),\end{split}$ (3)
[…]
Selective State Space Models (S6). Mamba [2] improves the performance of SSM by introducing Selective State Space Models (S6), allowing the continuous parameters to vary with the input enhances selective information processing across sequences, which extend the discretization process by selection mechanism:
$\bar{\bm{B}}$ $=s_{\bm{B}}(x),$ (4)
$\bar{\bm{C}}$ $=s_{\bm{C}}(x),$
$\Delta$ $=softplus\big(\text{Parameter}+s_{\bm{A}}(x)\big),$
where $s_{\bm{B}}(x)$ and $s_{\bm{C}}(x)$ are linear functions that project input $x$ into an $N$-dimensional space, while $s_{\bm{A}}(x)$ broadens a $D$-dimensional linear projection to the necessary dimensions. Parameter represents a trainable parameter matrix. softplussoftplus is a mathematical operation defined as: $softplus(x)=log(1+e^{x})$. In this paper, we use the term state space model (SSM) to refer to variants of SSMs, including S6.
Figure 1: The overview of the Vim model.
Overview of Vim. As mentioned earlier, the effectiveness of our work is demonstrated through the improvement of Vim. Therefore, we provide a brief introduction to the overall architecture of Vim. As shown in Fig. 1, Vim first reshapes the 2D image $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{n\times(p^{2}\cdot c)}$, where $c$ is the number of channels, $(h,w)$ is the resolution of the original image, and $(p,p)$ is the resolution of each image patch. The effective sequence length for Vim is therefore $n=hw/p^{2}$. Then, Vim linearly projects $\mathbf{X}_{p}$ into vectors of size $d$ and adds them to the position embeddings $E_{pos}\in\mathbb{R}^{(n+1)\times d}$, as follows:
@@
where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of blocks, and $\mathbf{Norm}$ is the normalization layer. The main difference between the Vim block and the classic Mamba block [2] resides in that Vim introduces a bidirectional scanning strategy, employing two independent branches to process the forward and backward sequences respectively. Each branch comprises a convolutional layer and an S6 module.
3 Rethinking the Scanning Strategy of ViMs
@@
The non-causal nature of two-dimensional spatial patterns in images is inherently inconsistent with the causal processing framework of SSM. Therefore, it is meaningful to explore methods for serializing images. However, most existing works on scanning strategies are typically based on a rather idealized assumption, that is, introducing more or more complex one-dimensional sequence scanning strategies may help enhance Mamba’s ability to understand local two-dimensional correlations in images. However, most existing works on scanning strategies are usually based on a rather idealized assumption. Scanning the same sequence multiple times in different orders can expand Mamba’s receptive field in different directions, thereby enhancing Mamba’s ability to understand local 2D correlations in images. Theoretically, regardless of the order, reducing the distance between adjacent tokens in one direction in two-dimensional space inevitably increases the distance between adjacent tokens in another direction, thus these strategies are essentially equivalent. Additionally, in the absence of global information serving as a reference, these sorting methods are essentially disconnected, making it inherently difficult for Mamba to gain additional spatial relationship information by scanning multiple sequences.
In practice, most works involving scanning strategy research, including Vim [22], LocalMamba [5], EfficientVMamba [14], and PlainMamba [20], lack clear comparative experiments to verify the necessity and effectiveness of their scanning strategies. VMamba [9] conducted experiments on different scanning strategies using ImageNet-1K, and the results indicate that most strategies show no significant performance differences. Zhu et al. [23] conducted a comprehensive experimental investigation on the impact of mainstream scanning directions and their combinations on semantic segmentation of remotely sensed images. Through extensive experiments on the LoveDA, ISPRS Potsdam, and ISPRS Vaihingen datasets, they demonstrate that no single scanning strategy outperforms others, regardless of their complexity or the number of scanning directions involved. In summary, an effective scanning strategy needs to maintain a global receptive field throughout the scanning process and demonstrate its effectiveness through analysis and comparative experiments.
4 Methodology
@@
In this section, we first elaborate on the advantages of frequency-domain scanning and the scanning strategy we proposed. Subsequently, we present the specific details of patch embedding in Vim-F. We then discuss the composition of the Vim-F and Vim-F(H) blocks. Finally, we explore the significance and feasibility of the convolution-free visual Mamba encoder and propose a proof-of-concept network, Vim-F(CF).
4.1 Scanning in the Frequency Domain
As analyzed in Section 3, maintaining a global receptive field during the scanning process is crucial for enhancing Mamba’s ability to understand local 2D correlations in images through improved scanning strategies. We believe that introducing frequency-domain scanning is a simple and straightforward solution. Specifically, we first perform a Fourier transform on the feature map to obtain:
$F(u,v)=\sum_{x=0}^{H-1}\sum_{y=0}^{W-1}f(x,y)e^{-j2\pi\left(\frac{ux}{H}+\frac{vy}{W}\right)},$ (7)
where $j$ represents the imaginary unit, $H$ and $W$ represent the height and width of the feature map, respectively. $f(x,y)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(x,y)$ in the feature map, and $F(u,v)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(u,v)$ in the frequency domain. From Eq. (7), it can be seen that each frequency-domain coefficient $F(u,v)$ is related to the values of all pixels in the feature map. $|F(u,v)|$ represents the modulus of $F(u,v)$, which is also known as the amplitude spectrum. Obviously, scanning on the amplitude spectrum can ensure a global receptive field throughout the scanning process.
In addition, the Fourier transform has translation invariance. Specifically, if a function is translated in the spatial domain, its amplitude spectrum is the same as that of the original function. A brief proof is provided below: According to Eq. (7), assuming the translated feature map is f(xa,yb)f(x-a,y-b), where $a$ and $b$ are the translation amounts. The Fourier transform of the translated feature map is:
$F^{\prime}(u,v)=\sum_{x=0}^{H-1}\sum_{y=0}^{W-1}f(x-a,y-b)e^{-j2\pi\left(\frac{ux}{H}+\frac{vy}{W}\right)}.$ (8)
Let $x^{\prime}=x-a$, $y^{\prime}=y-b$, then $x=x^{\prime}+a$, $y=y^{\prime}+b$. Substituting into Eq. (8), we get:
$F^{\prime}(u,v)$ $=\sum_{x^{\prime}}\sum_{y^{\prime}}f(x^{\prime},y^{\prime})e^{-j2\pi\left(\frac{u(x^{\prime}+a)}{H}+\frac{v(y^{\prime}+b)}{W}\right)}$ (9)
$=e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}\sum_{x^{\prime}}\sum_{y^{\prime}}f(x^{\prime},y^{\prime})e^{-j2\pi\left(\frac{ux^{\prime}}{H}+\frac{vy^{\prime}}{W}\right)}$
$=e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}F(u,v).$
We know that the modulus of the phase factor $e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}$ is 1, so the amplitude spectrum of $F^{\prime}(u,v)$ is:
$|F^{\prime}(u,v)|=|e^{-j2\pi\left(\frac{ua}{H}+\frac{vb}{W}\right)}F(u,v)|=|F(u,v)|.$ (10)
This property makes the amplitude spectrum an ideal “reference object” during scanning, so simultaneous scanning in both the frequency and spatial domains is better than scanning in either domain alone. As shown in Fig. 2, adding the amplitude spectrum to the original feature map introduces global information while reducing inductive biases from one-dimensional serialization, thereby enhancing Mamba’s ability to understand semantic objects.
Figure 2: Illustrating the scanning strategy of Vim-F through an example of forward scanning in both the frequency and spatial domains. Due to the translation invariance of the Fourier transform, the amplitude spectra of figures (a) and (b) are identical.
4.2 Patch Embedding